# Tensor Product Attention Is All You Need

### v3

@@ B.1 RoPE Compatibility in Higher-Order TPA
where $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$, $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$, with $d_{c}=\tfrac{d_{h}}{d_{b}}$. Define the RoPE-transformed query as $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)}=\mathbf{Q}_{t}\mathbf{T}_{t}$, where
$\mathbf{T}_{t}=\mathbf{R}_{t}\otimes\mathbf{I}_{d_{c}}=\begin{pmatrix}\mathbf{R}_{t}&\cdots&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{R}_{t}&\cdots&\mathbf{0}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&\mathbf{R}_{t}\\ \end{pmatrix}\in\mathbb{R}^{d_{h}\times d_{h}},$
and $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ ($d_{b}\in\mathbb{Z}_{+}$ is even) is a block-diagonal matrix composed of $2\times 2$ rotation matrices:
@@
for $t\in\{1,\dots,T\}$ and $j\in\{1,\dots,d_{b}/2\}$.
This construction ensures that RoPE rotates only the coordinates corresponding to $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ while leaving $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ unchanged. Under these conditions, the RoPE-transformed query $\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)}$ admits a higher-order TPA factorization of the same rank $R_{Q}$. Specifically, we have
$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\Bigl{(}\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\Bigr{)}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)},$ (B.1)
@@
Please see Appendix C.2 for the proof. For fourth-order or higher, this result still holds.
Appendix C Proofs of Theorems
C.1 Proof of Theorem 1
@@
This equality confirms that the relative positional encoding between queries and keys is preserved under TPA’s factorization and RoPE’s rotation. Thus, TPA maintains compatibility with RoPE. This completes the proof of Theorem 1. ∎
C.2 Proof of Theorem 2
@@
We begin by observing that each term $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}$ is an element of $\mathbb{R}^{h}\otimes\mathbb{R}^{d_{h}}$. Here, $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$, $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$, with $d_{c}=\tfrac{d_{h}}{d_{b}}$. Consequently, the tensor product $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ forms a $d_{b}\times d_{c}$ matrix, and its vectorization lies in $\mathbb{R}^{d_{b}\cdot d_{c}}=\mathbb{R}^{d_{h}}$.
Applying the RoPE transformation to a single summand yields
$\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}\mapsto\mathbf{T}_{t}\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}.$
Since $\mathbf{T}_{t}$ is defined as the Kronecker product $\mathbf{R}_{t}\otimes\mathbf{I}_{d_{c}}$, where $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ and $\mathbf{I}_{d_{c}}$ is the identity matrix of size $d_{c}\times d_{c}$, it follows that
$\mathbf{T}_{t}\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}=\operatorname{vec}\bigl{(}\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}.$
This is because the Kronecker product with an identity matrix effectively applies the rotation $\mathbf{R}_{t}$ to the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ component while leaving $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ unchanged.
Therefore, the RoPE transformation of a single summand becomes
$\operatorname{RoPE}_{t}\Bigl{(}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}\Bigr{)}=\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}.$
Importantly, this transformation does not mix the components $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ and $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$; it solely rotates $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ via $\mathbf{R}_{t}$.
Summing over all ranks $r=1,\dots,R_{Q}$, we obtain
$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)},$
which retains the same higher-order TPA structure with rank $R_{Q}$.
Thus, the RoPE transformation is fully compatible with higher-order TPA, preserving the factorization rank and maintaining the structure by only rotating the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ components while leaving $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ unchanged. ∎
Appendix D More Related Works
@@
Low-Rank Factorizations. Low-rank approximations have been applied to compress model parameters and reduce complexity including LoRA (Hu et al., 2022), which factorizes weight updates during fine-tuning, and its derivatives for other training scenarios such as efficient pretraining (ReLoRA (Lialin et al., 2023), MoRA (Jiang et al., 2024)), long-context training (LongLoRA (Chen et al., 2024), SinkLoRA (Zhang, 2024)), as well as continual training (InfLoRA (Liang & Li, 2024), GS-LoRA (Zhao et al., 2024), I-LoRA (Ren et al., 2024)). These approaches typically produce static low-rank expansions that do not explicitly depend on the input context. And Malladi et al. (2023); Zeng & Lee (2024) provided theoretical proof of the expressiveness of low-rank approximation. For the initialization of factorization matrices, OLoRA (Büyükakyüz, 2024) applied QR-decomposition of pretrained weight to achieve better performance of language models while LoLDU (Shi et al., 2024) used LDU-decomposition to accelerate training of LoRA. Moreover, AdaLoRA (Zhang et al., 2023a) utilized Singular Value Decomposition (SVD) of the pretrained weight and introduced importance score for each parameter as a measurement to achieve dynamic adjustment of rank. TPA, by contrast, constructs Q, K, and V as contextually factorized tensors, enabling dynamic adaptation.
Appendix E More on Attention Mechanisms
E.1 Multi-Query Attention (MQA)

### v4

@@ E.1 RoPE Compatibility in Higher-Order TPA
where $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$, $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$, with $d_{h}=d_{b}d_{c}$. Define the RoPE-transformed query as $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)}=\mathbf{Q}_{t}\mathbf{T}_{t}$, where
$\mathbf{T}_{t}=\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top}=\begin{pmatrix}(\mathbf{R}_{t})^{\top}&\cdots&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&(\mathbf{R}_{t})^{\top}&\cdots&\mathbf{0}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&(\mathbf{R}_{t})^{\top}\\ \end{pmatrix}\in\mathbb{R}^{d_{h}\times d_{h}},$
$\mathbf{I}_{d_{c}}$ is the identity matrix of size $d_{c}\times d_{c}$, and $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ ($d_{b}\in\mathbb{Z}_{+}$ is even) is the standard RoPE block-diagonal matrix composed of $2\times 2$ rotation matrices:
@@
for $t\in\{1,\dots,T\}$ and $j\in\{1,\dots,d_{b}/2\}$. The transformation $\mathbf{T}_{t}=\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top}$ operates on the $d_{h}$-dimensional vectorized features by post-multiplication. This structure of $\mathbf{T}_{t}$ ensures that the rotation effectively applied to the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ component (which is a column vector) corresponds to a pre-multiplication by $\mathbf{R}_{t}$, as detailed in the proof (Appendix F.2). This preserves the structure induced by $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ while rotating $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$.
Under these conditions, the RoPE-transformed query $\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)}$ admits a higher-order TPA factorization of the same rank $R_{Q}$:
$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\Bigl{(}\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\Bigr{)}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)},$ (E.1)
@@
Please see Appendix F.2 for the proof. For fourth-order or higher, this result still holds.
To assess its empirical performance, we implemented third-order TPA. Table 4 lists the evaluation results for a small model. These results provide an initial indication of its viability. A comprehensive comparison with second-order TPA variants of similar parameter counts or ranks would be necessary to fully evaluate the trade-offs.
Table 4: The evaluation results of small models with third-order TPA pre-trained using FineWeb-Edu 100B dataset with lm-evaluation-harness. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Few-shot ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
0-shot 49.24 24.91 57.06 34.01 31.80 63.33 50.59 23.23 66.90 44.56
2-shot 53.37 25.34 48.78 34.00 29.20 62.79 52.33 26.41 75.30 45.28
Table 5: The evaluation results of medium models with third-order TPA pre-trained using FineWeb-Edu 100B dataset with lm-evaluation-harness. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Few-shot ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
0-shot 56.19 29.44 58.36 43.74 33.80 67.95 52.80 24.07 77.20 49.28
2-shot 62.84 30.89 55.69 43.16 33.00 68.01 51.78 26.57 86.90 50.98
Appendix F Proofs of Theorems
F.1 Proof of Theorem 3.1
@@ Appendix F Proofs of Theorems
This equality confirms that the relative positional encoding between queries and keys is preserved under TPA’s factorization and RoPE’s rotation. Thus, TPA maintains compatibility with RoPE. This completes the proof of Theorem 3.1. ∎
F.2 Proof of Theorem E.1
Theorem E.1 addresses the compatibility of RoPE with higher-order (specifically, 3rd-order) Tensor Product Attention. The theorem considers the query factorization:
where $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{h}$ (column vector), $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$ (column vector), $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$ (column vector), and $d_{h}=d_{b}d_{c}$. The term $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ is interpreted as the matrix $\mathbf{M}_{r}=\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})(\mathbf{c}_{r}^{Q}(\mathbf{x}_{t}))^{\top}\in\mathbb{R}^{d_{b}\times d_{c}}$. The notation $\mathbf{a}\otimes\mathbf{v}$ for $\mathbf{a}\in\mathbb{R}^{h}$ and $\mathbf{v}\in\mathbb{R}^{d_{h}}$ (column vectors) implies the outer product $\mathbf{a}\mathbf{v}^{\top}$. Thus, $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})(\operatorname{vec}(\mathbf{M}_{r}))^{\top}$.
The RoPE-transformed query is defined as $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)}=\mathbf{Q}_{t}\mathbf{T}_{t}$. Crucially, for the theorem’s conclusion to hold as intended (i.e., that the $\mathbf{b}_{r}^{Q}$ component is transformed by pre-multiplication with the standard RoPE matrix $\mathbf{R}_{t}$), the global transformation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ (that post-multiplies $\mathbf{Q}_{t}$) is given by:
$\mathbf{T}_{t}=\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top},$
where $\mathbf{I}_{d_{c}}$ is the $d_{c}\times d_{c}$ identity matrix, and $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ is the standard RoPE block-diagonal matrix that pre-multiplies $d_{b}$-dimensional column vectors (as defined explicitly in the theorem statement in Section E).
The theorem claims that, under these conditions, $\widetilde{\mathbf{Q}}_{t}$ admits a higher-order TPA factorization:
$\widetilde{\mathbf{Q}}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\Bigl{(}\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\Bigr{)},$
@@
Let $\mathbf{a}_{r}^{Q}\equiv\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})$, $\mathbf{b}_{r}^{Q}\equiv\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$, and $\mathbf{c}_{r}^{Q}\equiv\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ for brevity. Let $\mathbf{M}_{r}=\mathbf{b}_{r}^{Q}(\mathbf{c}_{r}^{Q})^{\top}\in\mathbb{R}^{d_{b}\times d_{c}}$. Let $\mathbf{v}_{r}=\operatorname{vec}(\mathbf{M}_{r})\in\mathbb{R}^{d_{h}}$ be the column vector obtained by stacking the columns of $\mathbf{M}_{r}$. The query tensor is $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{v}_{r})^{\top}$.
The RoPE transformation is $\widetilde{\mathbf{Q}}_{t}=\mathbf{Q}_{t}\mathbf{T}_{t}$. Substituting the factorization and the revised definition of $\mathbf{T}_{t}$:
$\widetilde{\mathbf{Q}}_{t}$ $=\left(\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{v}_{r})^{\top}\right)(\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top})$
$=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}\left((\mathbf{v}_{r})^{\top}(\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top})\right).$
Let’s analyze the transformed vector part for the $r$-th component: $(\mathbf{v}_{r})^{\top}(\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top})$. This row vector is the transpose of $((\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top})^{\top}\mathbf{v}_{r})$. Let’s compute the pre-multiplying matrix:
$((\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top})^{\top}=(\mathbf{I}_{d_{c}})^{\top}\otimes((\mathbf{R}_{t})^{\top})^{\top}=\mathbf{I}_{d_{c}}\otimes\mathbf{R}_{t}.$
So, the column vector transformation is $(\mathbf{I}_{d_{c}}\otimes\mathbf{R}_{t})\mathbf{v}_{r}$. Substitute $\mathbf{v}_{r}=\operatorname{vec}(\mathbf{M}_{r})=\operatorname{vec}(\mathbf{b}_{r}^{Q}(\mathbf{c}_{r}^{Q})^{\top})$:
$(\mathbf{I}_{d_{c}}\otimes\mathbf{R}_{t})\operatorname{vec}(\mathbf{b}_{r}^{Q}(\mathbf{c}_{r}^{Q})^{\top}).$
We use the Kronecker product identity: $(\mathbf{B_{0}}^{\top}\otimes\mathbf{A_{0}})\operatorname{vec}(\mathbf{X_{0}})=\operatorname{vec}(\mathbf{A_{0}}\mathbf{X_{0}}\mathbf{B_{0}})$. To match our expression $(\mathbf{I}_{d_{c}}\otimes\mathbf{R}_{t})\operatorname{vec}(\mathbf{M}_{r})$, we identify: $\mathbf{A_{0}}=\mathbf{R}_{t}$, $\mathbf{B_{0}}^{\top}=\mathbf{I}_{d_{c}}\implies\mathbf{B_{0}}=\mathbf{I}_{d_{c}}$, $\mathbf{X_{0}}=\mathbf{M}_{r}=\mathbf{b}_{r}^{Q}(\mathbf{c}_{r}^{Q})^{\top}$. Applying the identity, we get:
$\operatorname{vec}\left(\mathbf{R}_{t}(\mathbf{b}_{r}^{Q}(\mathbf{c}_{r}^{Q})^{\top})\mathbf{I}_{d_{c}}\right)=\operatorname{vec}\left((\mathbf{R}_{t}\mathbf{b}_{r}^{Q})(\mathbf{c}_{r}^{Q})^{\top}\right).$
Let $\widetilde{\mathbf{b}}^{Q}_{r}=\mathbf{R}_{t}\mathbf{b}_{r}^{Q}$. This is precisely the transformation for the $\mathbf{b}_{r}^{Q}$ component as claimed in the theorem. So the transformed column vector is $\operatorname{vec}(\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{c}_{r}^{Q})^{\top})$. The corresponding row vector in the sum for $\widetilde{\mathbf{Q}}_{t}$ is therefore $(\operatorname{vec}(\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{c}_{r}^{Q})^{\top}))^{\top}$.
Substituting this back into the expression for $\widetilde{\mathbf{Q}}_{t}$:
$\widetilde{\mathbf{Q}}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\operatorname{vec}(\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{c}_{r}^{Q})^{\top}))^{\top}.$
This is equivalent to the theorem’s claimed factorization, using the definition $\mathbf{a}\otimes\mathbf{col\_vec}=\mathbf{a}(\mathbf{col\_vec})^{\top}$:
$\widetilde{\mathbf{Q}}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}\otimes\operatorname{vec}\bigl{(}\widetilde{\mathbf{b}}^{Q}_{r}\otimes\mathbf{c}_{r}^{Q}\bigr{)},$
where $\widetilde{\mathbf{b}}^{Q}_{r}=\mathbf{R}_{t}\mathbf{b}_{r}^{Q}$. This completes the proof, showing that RoPE can be consistently applied to higher-order TPA representations if the global RoPE transformation matrix $\mathbf{T}_{t}$ (that post-multiplies $\mathbf{Q}_{t}$) is appropriately defined as $\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top}$, ensuring that the standard RoPE matrix $\mathbf{R}_{t}$ effectively pre-multiplies the $\mathbf{b}_{r}^{Q}$ component. ∎
Appendix G More Related Work
Transformers and Attention. As a sequence-to-sequence architecture, Transformer [59] introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies. Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns [9, 48, 15, 29, 26, 30], kernel-based projections [10], and linearized transformers [58, 24, 43, 68, 53, 66]. To decrease memory usage and circumvent the limitation of memory bandwidth in training, [45] proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head. To tackle the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) [2] divides queries into several groups, and each group of queries shares a single key head and value head. Recently, DeepSeek-V2 [31] applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value. Concurrently, [20] proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q. Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
KV Cache Optimization. During the auto-regressive inference of Transformers, key and value (KV) tensors from previous tokens are cached to avoid recomputation, a technique first proposed by [39]. This Key-Value (KV) cache, while crucial for efficiency, consumes significant memory and can introduce latency bottlenecks due to memory bandwidth limitations [1]. Consequently, various studies have explored methods to mitigate these issues. These include KV cache eviction strategies that discard less significant tokens [69, 61, 7, 1], dynamic sparse attention mechanisms focusing on selected keys and values [41, 54, 49], offloading the KV cache to CPU memory [16, 25, 52], and quantizing the KV cache [60, 33, 18]. In contrast to these approaches, TPA focuses on reducing the intrinsic size of the KV cache by employing tensor-decomposed key and value representations.
@@ Appendix G More Related Work
Low-Rank Factorizations. Low-rank approximations are widely used to compress model parameters and reduce computational complexity. Notable examples include LoRA [19], which factorizes weight updates during fine-tuning, and its derivatives tailored for various training scenarios such as efficient pretraining (ReLoRA [27], MoRA [21]), long-context training (LongLoRA [8], SinkLoRA [65]), and continual training (InfLoRA [28], GS-LoRA [70], I-LoRA [40]). These methods generally produce static low-rank expansions that are independent of the input context. Theoretical justifications for the expressiveness of low-rank approximations have been provided by [37, 64]. Initialization strategies for these factorization matrices have also been explored: OLoRA [6] utilizes QR-decomposition of pretrained weights for improved language model performance, while LoLDU [47] employs LDU-decomposition to accelerate LoRA training. Furthermore, AdaLoRA [67] uses Singular Value Decomposition (SVD) on pretrained weights and introduces parameter importance scores to dynamically adjust ranks. TPA, in contrast, constructs Q, K, and V tensors using contextually-aware factorizations, allowing for dynamic adaptation based on the input.
Appendix H More on Attention Mechanisms
H.1 Multi-Query Attention (MQA)